The length of a rectangle is 3m more than twice its width, and the area of the rectangle is 77 square meters. Find the dimensions of the rectangle.
Answer by RAY100(1637)

Area = Length * Width = (2w+3) * (w) = 77
.
2w^2 +3w -77 = 0
,
(2w-11)(w+7) =0
.
2w-11 = 0,,,,w=11/2=5.5 m
.
w+7 =0,,,,,,,,w=-7 (not reasonable)
.
L= 2w+3 = 14 m
.
check,,,,,A = L*W = 14*5.5= 77,,,,ok